VIBRATIONALLY RESOLVED NEGATIVE ION PHOTOELECTRON SPECTROSCOPY OF Nb_{8}

\underline{T. P. MARCY} and D. G. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455..

We have observed a surprisingly simple, vibrationally resolved spectrum for the bare niobium octamer cluster by photoelectron spectroscopy of mass selected Nb_{8}^{−} anions. The clusters were prepared in a cw cathode discharge source within a liquid nitrogen cooled flow tube (0.5 Torr), and the spectra were obtained at 488 nm (2.540 eV) with an electron kinetic energy resolution of about 5 meV (40 cm^{−1}). The measured electron affinity of 1.513(7) eV is consistent with that obtained in the earlier anion PES study ^{a}. In addition, the present spectrum exhibits a single clearly resolved vibrational progression in the transition between the ground electronic states of the anionic and neutral clusters. The frequency of this active mode is 185(15) cm^{−1} in the neutral molecule and 160(20) cm^{−1} in the anion, and the short progression indicates a small normal mode displacement of about 0.5 amu^{1/2}Å. These results are discussed in the context of recently reported density functional calculations ^{b} on Nb8 and Nb8^{−}.

aH. Kietzmann, J. Morenzin, P. S. Bechthold, G. Ganteför, W. Eberhardt, J. Chem. Phys. 109, 2275 (1998).

bH. Grönbeck, A. Rosén and W. Andreoni, Phys. Rev. A 58, 4630 (1998); R. Fournier, T. Pang and C. Chen, Phys. Rev. A, 57, 3683 (1998).